# A General Framework for Joint Multi-State Models

### v3

@@ Abstract
We propose a general multi-state joint modeling framework that unifies longitudinal biomarker dynamics with multi-state time-to-event processes defined on arbitrary directed graphs. The proposed framework also accomodates nonlinear longitudinal submodels and scalable inference via stochastic gradient descent. This formulation encompasses both Markovian and semi-Markovian transition structures, allowing recurrent cycles and terminal absorptions to be naturally represented. The longitudinal and event processes are linked through shared latent structures within nonlinear mixed-effects models, extending classical joint modeling formulations.
We derive the complete likelihood, model selection criteria, and develop scalable inference procedures based on stochastic gradient descent to enable high-dimensional and large-scale applications. In addition, we formulate a dynamic prediction framework that provides individualized state-transition probabilities and personalized risk assessments along complex event trajectories.
@@ 1 Introduction
Joint modeling of longitudinal and time-to-event data has become an essential tool of modern biostatistics [papageorgiou2019overview], particularly for dynamic prediction in clinical applications. Classical joint models typically couple a longitudinal biomarker process with a single time-to-event outcome [rizopoulos2012joint], allowing for the integration of biological knowledge and individual heterogeneity via shared latent structures. However, many real-world processes involve multiple possible outcomes, intermediate stages, or recurrent events, which cannot be fully captured by a single-event framework [KrolMauguen2017]. To address these limitations, specialized frameworks have been developed to handle competing risks and recurrent events. However, these remain constrained to specific event types and fail to accommodate more general multi-state transitions. In such settings, multi-state models provide a more general and flexible approach [ferrer2016joint].
@@
The link between multi-state models and joint models arises when a multi-state process is integrated as a component within a broader joint modeling framework . While multi-state models, such as those described in the “sequential state framework” primarily focus on movements between discrete states, joint models operate under a “parallel trajectory framework” that combines a longitudinal process with an acyclic time-to-event process.
The model proposed by Ferrer et al. [ferrer2016joint] exemplifies this link by presenting a joint model for a longitudinal process (e.g., Prostate-Specific Antigen (PSA) measurements) and a multi-state process (e.g., clinical progressions in prostate cancer). These two sub-models are interconnected by shared random effects, allowing the model to account for the correlation between the continuous longitudinal biomarker trajectory and the discrete transitions between health states.
@@
This paper is organized as follows. Section 2 reviews the background on multi-state and joint modeling frameworks. Section 3.2 introduces the proposed general multi-state joint modeling framework, detailing the likelihood formulation and proposed model selection creteria. Section 4 describes the stochastic-gradient–based inference algorithm, while Section 5 presents the dynamic prediction framework. Section 6 reports simulation experiments that assess parameter recovery and convergence properties of the proposed model, whereas Section 7 applies the framework to data from the PAQUID cohort. Finally, Section 8 concludes with perspectives and future research directions.
@@ 2.2 Multi-State Semi-Markov Processes
In an MSMP, the process is defined by a directed graph $G=(V,E)$, where $V$ represents the set of states and $E$ the set of possible transitions. The transition intensities $\lambda_{k^{\prime}\mid k}(t\mid t_{0})$ from state $k$ to state $k^{\prime}$ at time $t$, given entry time $t_{0}$ are therefore assumed to be non-identically zero. The sojourn time in state $k$ is not restricted to an exponential distribution, allowing for more realistic modeling of the time spent in each state.
MSMPs have been widely applied in a range of disciplines. In reliability, they are used to modeling degradation and repair processes [limnios2001semi]. In biomedical studies, they have proven useful for analyzing illness-death models or disease progression with non-exponential transitions [commenges2006likelihood, fiocco2008modelling]. Applications in finance also exist, where semi-Markovian dynamics can model credit rating migrations or economic regimes [luciano2006semi]. In these contexts, MSMPs retain the Markov property in the embedded jump chain while offering increased realism through flexible dwell time modeling.
@@ 2.3 Joint Models and Multi-State Processes
Although Ferrer et al. [ferrer2016joint] note that their joint modeling framework could be adapted to semi-Markov processes in other contexts, their primary application to prostate cancer progression assumes a non-homogeneous Markov multi-state process. However, this framework is not well suited for arbitrary transition patterns other than the ones modeled by DAGs.
While both joint models and multi-state models offer powerful tools for analyzing disease progression, they emphasize different aspects of the underlying process. Multi-state models capture the sequence and timing of discrete transitions between states, whereas joint models focus on the evolution of continuous longitudinal markers and their association with event risk. Existing joint multi-state models, such as that of Ferrer et al. [ferrer2016joint], are restricted to linear mixed-effects submodels and to acyclic Markov transition graphs, as their likelihood relies on a sequential factorization that is not valid when transitions form cycles. However, many biological processes involve cyclic or semi-Markov transitions and nonlinear biomarker dynamics. These cannot be handled by existing frameworks, motivating the present general formulation.
@@ 3 A General Multi-State Joint Modeling Framework
The proposed model defines a joint process $\{Y_{i}(t),S_{i}(t)\}_{t\geq 0}$, where $Y_{i}(t)$ denotes the longitudinal biomarker trajectory and $S_{i}(t)$ the multi-state event process. Both components are linked through shared latent structures, and the event process is represented on an arbitrary directed graph supporting both Markovian and semi-Markovian transition mechanisms.
@@ 3.1 Notation Conventions
Throughout this section, $i=1,\dots,n$ indexes individuals, and $t_{ij}$ denotes the $j$-th observation time of individual $i$. The vector $Y_{ij}\in\mathbb{R}^{d}$ represents longitudinal biomarker measurements, and $X_{i}\in\mathbb{R}^{p}$ the associated covariates.
The latent random effects $b_{i}\sim\mathcal{N}(0,Q)$ capture individual heterogeneity, while $\psi_{i}\coloneqq f(\gamma,X_{i},b_{i})$ denotes the subject-specific parameters in the longitudinal submodel. The filtration $\mathcal{H}_{i}(t)$ represents the observed history of the longitudinal process up to time $t$.
Transition times and states are denoted by $(T_{i\ell},S_{i\ell})$, with right-censoring time $C_{i}$. We also define the index of the last observed transition before time $t$ denoted by $m_{i}(t)\coloneqq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$.
3.2 Extending Multi-State Joint Models
The joint model proposed by Ferrer et al. [ferrer2016joint] established a foundational link between longitudinal biomarkers and multi-state processes through shared random effects. However, their formalism relies on binary censoring indicators to characterize transitions, thereby inherently restricting the modeling framework to Directed Acyclic Graphs (DAGs) and excluding recurrent events. To overcome this limitation, we present a new method based on arbitrary graph transitions. By shifting away from binary indicators toward a more flexible graph-theoretic specification, our approach accommodates complex disease trajectories, including recurrent events and cyclic transitions that cannot be captured within a DAG-based structure.
@@ 3.2.1 Graph Structure
Let $G=(V,E)$ be a directed graph, where $V=\{1,\dots,p\}$ denotes the set of states and $E\subset V\times V$ the set of admissible transitions. The graph encodes all possible paths of the multi-state process, allowing for competing, recurrent, or absorbing transitions. Figure 1 illustrates an example of a four-state model including an absorbing state (Death). The corresponding adjacency matrix $A=(A_{kk^{\prime}})_{1\leq k,k^{\prime}\leq|V|}$ satisfies $A_{kk^{\prime}}=1$ if $(k,k^{\prime})\in E$ and 0 otherwise.
@@ 3.2.2 Individual Trajectories
Each individual $i$ follows a latent trajectory
$\mathcal{T}_{i}^{*}\coloneqq\big((T_{i0},S_{i0}),(T_{i1},S_{i1}),(T_{i2},S_{i2}),\dots\big),$
where $T_{i\ell}$ denotes the $\ell$-th transition time and $S_{i\ell}\in V$ the corresponding state. The observed trajectory is right-censored at time $C_{i}$, so that only $\{(T_{i\ell},S_{i\ell}):T_{i\ell}\leq C_{i}\}$ is observed. Between transitions, the state $S_{i}(t)$ is piecewise constant. We note $m_{i}(t)\coloneq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$ the index of the last observed transition before time $t$ and $m_{i}=m_{i}(C_{i})$.
@@
$\mathscr{L}\big((T_{i,\ell+1},S_{i,\ell+1})\mid\{(T_{i\ell^{\prime}},S_{i\ell^{\prime}})\}_{\ell^{\prime}\leq\ell}\big)=\mathscr{L}\big((T_{i,\ell+1},S_{i,\ell+1})\mid(T_{i\ell},S_{i\ell})\big),$ (1)
@@ 3.2.3 Transition Intensities and Survival Functions
For any admissible transition $(k,k^{\prime})\in E$, let $\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})$ denote the instantaneous risk of moving from $k$ to $k^{\prime}$ at time $t$, given entry time $t_{0}$
$\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})\coloneqq\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i,\ell+1}\leq t+\delta,S_{i,\ell+1}=k^{\prime}\mid T_{i,\ell+1}>t,T_{i\ell}=t_{0},S_{i\ell}=k)}{\delta}.$
The cumulative risk and the corresponding survival function are
$\Lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})=\int_{t_{0}}^{t}\lambda_{i}^{k^{\prime}\mid k}(w\mid t_{0})\,dw,$
$\mathbb{P}(T_{i,\ell+1}>t\mid T_{i\ell},S_{i\ell})=\exp\left[-\sum_{s:(S_{i\ell},s)\in E}\Lambda_{i}^{s|S_{i\ell}}(t\mid T_{i\ell})\right].$
@@
$\mathbb{P}(S_{i,\ell+1}=k^{\prime}\mid T_{i,\ell+1},T_{i\ell},S_{i\ell})=\frac{\lambda_{i}^{k^{\prime}|S_{i\ell}}(T_{i,\ell+1}\mid T_{i\ell})}{\sum_{s:(S_{i\ell},s)\in E}\lambda_{i}^{s|S_{i\ell}}(T_{i,\ell+1}\mid T_{i\ell})}.$
@@ 3.3 Multi-State Joint Model Specification
The proposed multi-state joint model with Gaussian prior and homoscedastic Gaussian noise is specified by:
$\begin{cases}Y_{ij}=h(t_{ij},\psi_{i})+\epsilon_{ij},\quad\epsilon_{ij}\sim\mathcal{N}(0,R),\\ \psi_{i}=f(\gamma,X_{i},b_{i}),\quad b_{i}\sim\mathcal{N}(0,Q),\\ \lambda_{i}^{k^{\prime}\mid k}\left(t\mid X_{i},T_{im_{i}(t)},\mathcal{H}_{i}(t)\right)=\lambda_{0}^{k^{\prime}\mid k}(t\mid T_{im_{i}(t)})\exp\big(\alpha^{k^{\prime}\mid k}g^{k^{\prime}\mid k}(t,X_{i},\psi_{i})+\beta^{k^{\prime}\mid k}X_{i}\big).\end{cases}$
Here, $h$ and $f$ define the nonlinear mixed-effects submodel, $g^{k^{\prime}\mid k}$ represents the link between biomarker dynamics and the transition intensity, and $\lambda_{0}^{k^{\prime}\mid k}$ denotes the baseline hazard associated with transition $(k\to k^{\prime})$. This structure extends the classical joint model to arbitrary directed graphs and to both Markovian and semi-Markov specifications.
The coefficient $\alpha^{k^{\prime}\mid k}$ quantifies the effect of the longitudinal biomarker dynamics on the instantaneous risk of transitioning from state $k$ to state $k^{\prime}$. More concretely, for a one-unit increase in $g^{k^{\prime}\mid k}(t,X_{i},\psi_{i})$ (holding other covariates constant), the hazard ratio for transition $(k\to k^{\prime})$ is $\exp(\alpha^{k^{\prime}\mid k})$. Likewise, $\beta^{k^{\prime}\mid k}$ represents the effect of baseline (or time-varying) covariates $X_{i}$ on that same transition: a one-unit increase in a covariate $X_{ij}$ multiplies the hazard by $\exp(\beta^{k^{\prime}\mid k}_{j})$. Thus, $\alpha^{k^{\prime}\mid k}$ captures the degree to which the biomarker trajectory (via its latent parameter $\psi_{i}$ or function $g^{k^{\prime}\mid k}$) influences the transition risk, while $\beta^{k^{\prime}\mid k}$ captures the direct effect of covariates on that transition’s risk, beyond the biomarker pathway.

### v4

@@ Abstract
We propose a general multi-state joint modeling framework that unifies longitudinal biomarker dynamics with multi-state time-to-event processes defined on arbitrary directed graphs. The proposed framework also accommodates nonlinear longitudinal submodels and scalable inference via stochastic gradient ascent. This formulation encompasses both Markovian and semi-Markovian transition structures, allowing recurrent cycles and terminal absorptions to be naturally represented. The longitudinal and event processes are linked through shared latent structures within nonlinear mixed-effects models, extending classical joint modeling formulations.
We derive the complete likelihood, model selection criteria, and develop scalable inference procedures based on stochastic gradient ascent to enable high-dimensional and large-scale applications. In addition, we formulate a dynamic prediction framework that provides individualized state-transition probabilities and personalized risk assessments along complex event trajectories.
@@ 1 Introduction
Joint modeling of longitudinal and time-to-event data has become an essential tool of modern biostatistics [papageorgiou2019overview], particularly for dynamic prediction in clinical applications. Classical joint models typically couple a longitudinal biomarker process with a single time-to-event outcome [rizopoulos2012joint], allowing for the integration of biological knowledge and individual heterogeneity via shared latent structures. However, many real-world processes involve multiple possible outcomes, intermediate stages, or recurrent events, which cannot be fully captured by a single-event framework [KrolMauguen2017]. To address these limitations, specialized frameworks have been developed to handle competing risks and recurrent events. However, these remain constrained to specific event types and fail to accommodate more general multi-state transitions. In such settings, multi-state models provide a more general and flexible approach [ferrer2016joint, you2024joint].
@@
The link between multi-state models and joint models arises when a multi-state process is integrated as a component within a broader joint modeling framework. While multi-state models, such as those described in the “sequential state framework” primarily focus on movements between discrete states, joint models operate under a “parallel trajectory framework” that combines a longitudinal process with a time-to-event process.
The model proposed by Ferrer et al. [ferrer2016joint] exemplifies this link by presenting a joint model for a longitudinal process (e.g., Prostate-Specific Antigen (PSA) measurements) and a multi-state process (e.g., clinical progressions in prostate cancer). These two sub-models are interconnected by shared random effects, allowing the model to account for the correlation between a continuous longitudinal biomarker trajectory and acyclic discrete transitions between health states. Specifically, the previously proposed mechanism [ferrer2016joint, you2024joint] is based on binary censoring indicators for each state, implying that any given state can be entered at most once, therefore precluding the modeling of recurrent events.
@@
This paper is organized as follows. Section 2 reviews the background on multi-state and joint modeling frameworks. Section 3.2 introduces the proposed general multi-state joint modeling framework, detailing the likelihood formulation and proposed model selection criteria. Section 4 describes the stochastic-gradient–based inference algorithm, while Section 5 presents the dynamic prediction framework. Section 6 reports simulation experiments that assess parameter recovery and convergence properties of the proposed model, whereas Section 7 applies the framework to data from the PAQUID cohort [Letenneur1994]. Finally, Section 8 concludes with perspectives and future research directions.
@@ 2.2 Multi-State Semi-Markov Processes
In an MSMP, the process is defined by a directed graph $G=(V,E)$, where $V$ represents the set of states and $E$ the set of possible transitions. The transition intensities $\lambda_{k\to k^{\prime}}(t\mid t_{0})$ from state $k$ to state $k^{\prime}$ at time $t$, given entry time $t_{0}$, are therefore assumed to be non identically zero. The sojourn time in state $k$ is not restricted to an exponential distribution, allowing for more realistic modeling of the time spent in each state.
MSMPs have been widely applied in a range of disciplines. In reliability, they are used to model degradation and repair processes [limnios2001semi]. In biomedical studies, they have proven useful for analyzing illness-death models or disease progression with non-exponential transitions [commenges2006likelihood, fiocco2008modelling]. Applications in finance also exist, where semi-Markovian dynamics can model credit rating migrations or economic regimes [luciano2006semi]. In these contexts, MSMPs retain the Markov property in the embedded jump chain while offering increased realism through flexible dwell time modeling.
@@ 2.3 Joint Models and Multi-State Processes
While both joint models and multi-state models offer powerful tools for analyzing disease progression, they emphasize different aspects of the underlying process. Multi-state models capture the sequence and timing of discrete transitions between states, whereas joint models focus on the evolution of continuous longitudinal markers and their association with event risk. Existing joint multi-state models, such as that of Ferrer et al. [ferrer2016joint], are restricted to linear mixed-effects submodels and to acyclic Markov transition graphs, since their reliance on binary censoring indicators inherently prevents modeling recurrent events. However, many biological processes involve cyclic or semi-Markov transitions and nonlinear biomarker dynamics. These cannot be handled by existing frameworks, motivating the present general formulation.
@@ 3 A General Multi-State Joint Modeling Framework
The proposed model defines a joint process $\{Y_{i}(t),S_{i}(t)\}_{t}$, where $Y_{i}(t)$ denotes the longitudinal biomarker trajectory and $S_{i}(t)$ the multi-state event process. Both components are linked through shared latent structures, and the event process is represented on an arbitrary directed graph supporting both Markovian and semi-Markovian transition mechanisms.
@@ 3.1 Notation Conventions
Throughout this section, let $i\in\llbracket n\rrbracket$ index $n\geq 1$ individuals, and let $t_{ij}$ denote the $j$-th observation time of individual $i$. The vector $Y_{ij}\in\mathbb{R}^{d}$ represents longitudinal biomarker measurements, and $X_{i}\in\mathbb{R}^{p}$ the associated covariates.
The latent random effects $b_{i}\sim\mathcal{N}(0,Q)$ capture individual heterogeneity, $\gamma$ represent shared characteristics across all individuals, and the subject-specific parameters are defined as $\psi_{i}\coloneqq f(\gamma,X_{i},b_{i})$. The set $\mathcal{H}_{i}(t)\coloneqq\left\{(\sigma,h(s,\psi_{i})):s\leq t\right\}$ represents the true underlying history of the longitudinal process up to time $t$. Correspondingly, $\mathcal{Y}_{i}(t)$ denotes the observed noisy longitudinal history, and $\mathcal{S}_{i}(t)$ denotes the observed trajectory of the semi-Markovian dynamics.
Transition times and states are denoted by $(T_{i\ell},S_{i\ell})\in\mathbb{R}^{+}\times V$, where $V$ is the space of admissible states, and are subject to right-censoring at time $C_{i}\in\bar{\mathbb{R}}^{+}$. We define the index of the last observed transition before time $t$ as $m_{i}(t)\coloneqq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$. The complete sequence of transition times and states for individual $i$ is denoted by $\mathcal{T}_{i}^{*}$, whereas $\mathcal{T}_{i}$ denotes the corresponding right-censored sequence.
3.2 Multi-State Processes on Arbitrary Graphs
The joint model proposed by Ferrer et al. [ferrer2016joint] established a foundational link between longitudinal biomarkers and multi-state processes through shared random effects. However, their formalism relies on binary censoring indicators to characterize transitions, thereby inherently restricting the modeling framework to Directed Acyclic Graphs (DAGs) and excluding recurrent events. Indeed, the proposed model uses, for each state $k\in E$, a tuple $(T_{k},\delta_{k})$, where $T_{k}$ denotes the observed or censored time until leaving state $k$ and $\delta_{k}$ is an indicator equal to $1$ if the transition was observed and 0 if the observation was censored (i.e., the transition was not observed).
To overcome this limitation, we present a new method based on arbitrary graph transitions. By shifting away from binary indicators toward a more flexible graph-theoretic specification, our approach accommodates complex disease trajectories, including recurrent events and cyclic transitions that cannot be captured within a DAG-based structure.
@@ 3.2.1 Graph Structure
Let $G=(V,E)$ be a directed graph, where $V=\llbracket p\rrbracket$ denotes the set of states and $E\subseteq V\times V$ the set of admissible transitions. The graph encodes all possible paths of the multi-state process, allowing for competing, recurrent, or absorbing transitions. Figure 1 illustrates an example of a four-state model, including an absorbing state (Death). The corresponding adjacency matrix $A=(A_{kk^{\prime}})_{1\leq k,k^{\prime}\leq p}$ satisfies $A_{kk^{\prime}}=1$ if $(k,k^{\prime})\in E$ and 0 otherwise.
@@ 3.2.2 Individual Trajectories
Each individual $i\in\llbracket n\rrbracket$ follows a latent trajectory
$\mathcal{T}_{i}^{*}\coloneqq\left((T_{i0},S_{i0}),(T_{i1},S_{i1}),(T_{i2},S_{i2}),\dots\right),$
where $T_{i\ell}\in\mathbb{R}^{+}$ denotes the $\ell$-th transition time and $S_{i\ell}\in V$ the corresponding state. The observed trajectory is right-censored at time $C_{i}\in\bar{\mathbb{R}}^{+}$, so that only $\mathcal{T}_{i}\coloneqq\left\{(T_{i\ell},S_{i\ell}):T_{i\ell}\leq C_{i}\right\}$ is observed. Between transitions, the state $S_{i}(t)$ is constant, thus the transitions are the discontinuity points of the map $t\mapsto\mathcal{S}_{i}(t)$, of which we assume there are at most only infinitely countably many.
Letting $m_{i}(t)\coloneqq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$ be the index of the last observed transition before time $t$ and $m_{i}\coloneqq m_{i}(C_{i})$, we thus assume that no further transitions occur for individual $i$ on the interval $(T_{im_{i}},C_{i}]$. We also define $\mathcal{S}_{i}(t)\coloneqq\left\{(s,S_{i}(s)):s\leq t\right\}$ as the observed trajectory up to time $t$, which is equivalent to specifying all transition times and their corresponding states up to $t$, together with the time $t$ itself.
We emphasize once more that modeling the observed data solely as censored collections of transition times, as in Ferrer et al. [ferrer2016joint], is not sufficient to represent such trajectories, since it implicitly restricts each state to be visited at most once and therefore rules out recurrent transitions.
@@
$\mathscr{L}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid\{(T_{i\ell^{\prime}},S_{i\ell^{\prime}})\}_{\ell^{\prime}\leq\ell}\right)=\mathscr{L}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid(T_{i\ell},S_{i\ell})\right),$ (1)
@@ 3.2.3 Transition Intensities and Survival Functions
Although one could equivalently define intensities for all ordered pairs of states and set those corresponding to inadmissible transitions identically to zero, explicitly restricting attention to a graph of admissible transitions provides a clearer and more coherent representation of the underlying state dynamics.
For any admissible transition $(k,k^{\prime})\in E$ with nonzero probability, let $\lambda_{i}^{k\to k^{\prime}}(t_{0},t)$ denote the instantaneous risk at time $t$, given entry time $t_{0}\leq t$, and conditionally the transition $k\to k^{\prime}$
$\lambda_{i}^{k\to k^{\prime}}(t_{0},t)\coloneqq\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i(\ell+1)}\leq t+\delta\mid T_{i(\ell+1)}>t,T_{i\ell}=t_{0},S_{i\ell}=k,S_{i(\ell+1)}=k^{\prime})}{\delta}.$
Assuming this quantity to be finite for any pair $(t_{0},t)$ so that the underlying distribution of transition times is continuous, the cumulative risk and the corresponding survival function are
$\Lambda_{i}^{k\to k^{\prime}}(t_{0},t)=\int_{t_{0}}^{t}\lambda_{i}^{k\to k^{\prime}}(t_{0},s)\,ds,$
$\forall t\geq T_{i\ell},\,\mathbb{P}(T_{i(\ell+1)}>t\mid T_{i\ell},S_{i\ell})=\exp\left(-\sum_{j:(S_{i\ell},j)\in E}\Lambda_{i}^{S_{i\ell}\to j}(T_{i\ell},t)\right).$
@@
$\mathbb{P}(S_{i(\ell+1)}=k^{\prime}\mid T_{i(\ell+1)},T_{i\ell},S_{i\ell})=\frac{\lambda_{i}^{S_{i\ell}\to k^{\prime}}(T_{i\ell},T_{i(\ell+1)})}{\sum_{j:(S_{i\ell},j)\in E}\lambda_{i}^{S_{i\ell}\to j}(T_{i\ell},T_{i(\ell+1)})},$
so the joint density of a transition writes as
$p\left((T_{i(\ell+1)},T_{i(\ell+1)})\mid(T_{i\ell},T_{i\ell})\right)=$ $\frac{\lambda_{i}^{S_{i\ell}\to S_{i(\ell+1)}}(T_{i\ell},T_{i(\ell+1)})}{\sum_{j:(S_{i\ell},j)\in E}\lambda_{i}^{S_{i\ell}\to j}(T_{i\ell},T_{i(\ell+1)})}$
$\exp\left(-\sum_{j:(S_{i\ell},j)\in E}\Lambda_{i}^{S_{i\ell}\to j}(T_{i\ell},T_{i(\ell+1)})\right)$ (2)
In practice, numerical integration can be efficiently carried out using quadrature methods such as the Gauss-Legendre quadrature [lether1978construction].
@@ 3.3 Multi-State Joint Model Specification
Omitting the implicit conditioning on the individual parameters and covariates for the sake of simple notation, the proposed multi-state joint model with a Gaussian prior and homoscedastic Gaussian noise is specified by
$\begin{cases}Y_{ij}=h(t_{ij},\psi_{i})+\epsilon_{ij},\quad\epsilon_{ij}\sim\mathcal{N}(0,R),\\ \lambda_{i}^{k\to k^{\prime}}(t_{0},t)=\lambda_{0}^{k\to k^{\prime}}(t_{0},t^{\prime})\exp\left(\alpha^{k\to k^{\prime}}g^{k\to k^{\prime}}(t,X_{i},\psi_{i})+\beta^{k\to k^{\prime}}X_{i}\right),\\ \psi_{i}=f(\gamma,X_{i},b_{i}),\quad b_{i}\sim\mathcal{N}(0,Q).\end{cases}$
Here, $h$ and $f$ define the nonlinear mixed-effects submodel, $g^{k\to k^{\prime}}$ represents the link between biomarker dynamics and the transition intensity, and $\lambda_{0}^{k\to k^{\prime}}$ denotes the baseline hazard associated with transition $(k\to k^{\prime})$. This structure extends the classical joint model to arbitrary directed graphs and to both Markovian and semi-Markov specifications.
The coefficient $\alpha^{k\to k^{\prime}}$ quantifies the effect of the longitudinal biomarker dynamics on the instantaneous risk of transitioning from state $k$ to state $k^{\prime}$. More concretely, for a one-unit increase in $g^{k\to k^{\prime}}(t,X_{i},\psi_{i})$ (holding other covariates constant), the hazard ratio for transition $(k\to k^{\prime})$ is $\exp(\alpha^{k\to k^{\prime}})$. Likewise, $\beta^{k\to k^{\prime}}$ represents the effect of baseline (or time-varying) covariates $X_{i}$ on that same transition: a one-unit increase in a covariate $X_{ij}$ multiplies the hazard by $\exp(\beta^{k\to k^{\prime}}_{j})$. Thus, $\alpha^{k\to k^{\prime}}$ captures the degree to which the biomarker trajectory (via its latent parameters $\psi_{i}$ or function $g^{k\to k^{\prime}}$) influences the transition risk, while $\beta^{k\to k^{\prime}}$ captures the direct effect of covariates on that transition’s risk, beyond the biomarker pathway.